OpenStudy (wintersuntime):

Graph the following absolute value functions and write the corresponding piecewise functions for each. g(x)=|x^2-4|+1 Piecewise:

OpenStudy (wintersuntime):

Can someone please help me graph this and find the piecewise function

OpenStudy (anonymous):

this is going to be confusing

OpenStudy (wintersuntime):

yes

OpenStudy (anonymous):

the absolute value of \(\spadesuit\) is \(\spadesuit\) if \(\spadesuit>0\) otherwise it is \(-\spadesuit\)

OpenStudy (wintersuntime):

of what ?

OpenStudy (anonymous):

that is a spade

OpenStudy (wintersuntime):

ok

OpenStudy (anonymous):

lets write is as a piecewise function

OpenStudy (wintersuntime):

alright

OpenStudy (anonymous):

\[f(x) = |x| = \left\{\begin{array}{rcc} -x & \text{if} & x<0 \\ x& \text{if} & x\ge0 \end{array} \right.\]

OpenStudy (wintersuntime):

okay

OpenStudy (anonymous):

if x is negative, then the absolute value of \(x\) is \(-x\) (which is positive) and if \(x\) is positive then the absolute value of \(x\) is just \(x\)

OpenStudy (anonymous):

no you don't have \(x\) you have \(x^2-4\) right?

OpenStudy (wintersuntime):

oh okay

OpenStudy (wintersuntime):

yes

OpenStudy (anonymous):

so lets replace \(x\) in the above definition \[f(x) = |x| = \left\{\begin{array}{rcc} -x & \text{if} & x<0 \\ x& \text{if} & x\ge0 \end{array} \right.\] by \(x^2-4\)

OpenStudy (anonymous):

it is easy for me to do, i will just cut and paste

OpenStudy (wintersuntime):

okay

OpenStudy (anonymous):

\[f(x) = |x^2-4| = \left\{\begin{array}{rcc} -(x^2-4) & \text{if} & x^2-4<0 \\ x^2-4& \text{if} & x^2-4\ge0 \end{array} \right.\]

OpenStudy (anonymous):

we are clearly not done yet, but almost

OpenStudy (wintersuntime):

alright

OpenStudy (anonymous):

we just have to write the two inequalities above in terms of \(x\) instead of \(x^2-4\) first is it clear what i did? i literally copied and pasted \(x^2-4\) in to each \(x\)

OpenStudy (wintersuntime):

Yeah I see what you did

OpenStudy (anonymous):

ok good now how to we solve \[x^2-4\geq 0\]? do you know?

OpenStudy (wintersuntime):

Do we plug in numbers for x?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\(y=x^2-4\) is a parabola that opens up it is zero if \(x=-2\) or if \(x=2\)

OpenStudy (wintersuntime):

why -2 or 2 ?

OpenStudy (anonymous):